OpenStudy (dtan5457):

If one root of a quadratic equation is 3 and the product of the roots is 15? What is the value of b if the equation is in the form ax^2+bx+c=0? And a=1?

OpenStudy (dtan5457):

Is this 8? I just attempted to do it again.

Directrix (directrix):

sum of roots: −b/ a product of roots: c/a where the quadratic is in the form ax^2 + bx + c http://www.mathwarehouse.com/quadratic/roots/formula-sum-product-of-roots.php

OpenStudy (dtan5457):

I got 8 because if the product is 15, then c must be 15. therefore if one root is 3, then the other must be 5. so b=8??

Directrix (directrix):

I'm getting b = -8. I may have an error. I agree that c = 15. I'll post my other work.

Directrix (directrix):

a = 1, 3 is a given root 1* (3)^2 + b(3) + 15 = 0

Directrix (directrix):

@9 + 3b + 15 = 0 3b = -15 -9 3b = -24 b = -8 What do you think? @dtan5457

OpenStudy (dtan5457):

Not sure. I didn't change the value of a when they told me it equals one.